OpenStudy (anonymous):

In a triangle, the measure of the second angle is twice the measure of the first angle. The third angle is equal to the sum of the other angles.

OpenStudy (anonymous):

 because its a triangle, A + B + C = 180

OpenStudy (anonymous):

x + 2x + (x +2x) = Sum well known in all triangs ?

OpenStudy (anonymous):

from the information, B = 2A , C = A + 2A = 3A

OpenStudy (anonymous):

@jewjewbri simplify and solve

OpenStudy (anonymous):

therefore A + 2A + 3A = 180 (as @Mikael has already posted)

OpenStudy (anonymous):

Simplify and solve the equation you gave?

OpenStudy (anonymous):

yes but with 180 as RHS